Jeff starts driving at 55 mph from the same point that Lauren starts driving from 70 mph. They drive in opp. Directions and Lauren has a half hour head start. How long will they be able to talk on their cell phone and have a 440 mile range.
Answer by checkley75(3666)

55T+70(T+.5)=440
55T+70T+35=440
125T=440-35
125T=405
T=405/125
T=3.24 HOURS AFTER LAUREN LEAVES THEY WILL BE 440 MILES APART